P1 = 32 − 2Q1



P2 = 22 − 0.5Q2

1
Expert's answer
2021-12-28T08:41:57-0500

"\begin{cases}
 P=32-2Q, \\
 P=22-0.5Q,
\end{cases}"

"32-2Q=22-0.5Q,"

"10=1.5Q,"

"Q_1=Q_2=6.67,"

"P_1=P_2=18.67."